one is twice another number.When the larger number is diminished by 10, the result is two greater than the smaller. Find the numbers.
Answer by ewatrrr(24785)


Hi
one is twice another number
Let x and 2x represent the smaller number and larger number respectively
Question states***
2x - 10 = x+2
Solving for x
x = 12, the smaller number. The larger number is 24.
CHECKING our Answer***
24-10 = 12 + 2 = 14